OpenStudy (anonymous):

Which of the following is the equation of a line perpendicular to 12y = 3x + 5 that passes through (-1, 1)

OpenStudy (jmark):

First find the slope of the equation 12y = 3x+5 divide the equation by 12. y =( 3/12)x+5/12 slope m = 3/12 = ¼ the line perpendicular will have slope m = -4 equation of the line passing through the point ( -1,1) (y-y1) = m(x-x1) (y-1) = -4 (x-(-1) ) (y-1) = -4(x+1) y-1 = -4x-4 y = -4x-4+1 y = -4x-3 ICSE Results 2014 - http://icse.edurite.com/icse-isc-results-2014.html